1/3

Christian Morgenstern, Christian Morgenstern, Christian Morgenstern, Christian Morgenstern,

The Werewolf La lupfanto- mo Der Werwolf The Banshee (An Approach)

translated by Alexander Gross translated by Bertilo Wennergren translated by Max Knight

A Werewolf, troubled by

his name, Left wife and brood one

night and came To a hidden graveyard to

enlist The aid of a long-dead phi-

lologist.

Jen lupfantom’ de

famili’ meznokte iam ku-

ris for al tombo de la

profesor’. kaj petis pri kon-

jugaci’.

Ein Werwolf eines

Nachts entwich von Weib und Kind

und sich begab an eines Dorfschul-

lehrers Grab und bat ihn: �Bitte,

beuge mich!“ One night, a banshee slunk away from mate and child, and in the gloom went to a village teacher’s tomb, requesting him: “Inflect

me, pray.”

“Oh sage, wake up, please

don’t berate me,” He howled sadly, “Just con-

jugate me.” The seer arose a bit unstea-

dy Yawned twice, wheezed on-

ce, and then was ready.

Do diktis tiu dok-

ta ul’ starante sur la

kructabul’ dum lup’ senmova

kiel splinto atentis vortojn de

l’ mortinto.

Der Dorfschulmeister

stieg hinauf auf seines Blech-

schilds Messingknauf

und sprach zum Wolf,

der seine Pfoten geduldig kreuzte vor

dem Toten:

The village teacher climbed up straight upon his grave stone with its plate and to the apparition

said who meekly knelt before the dead:

“Well, ‘Werewolf’ is your

plural past, While ‘Waswolf’ is singu-

larly cast: There’s ‘Amwolf’ too, the

present tense, And ‘Iswolf,’ ‘Arewolf’ in

this same sense.”

De “lupfantomo”,

simpla rekta, tra “lupfintomo”,

plej perfekta, plu “lupfontomo”,

iom rara, ĝis “lupfuntomo”,

ho, koŝmara.

�Der Werwolf“ – sprach der gute Mann,

�des Werwolfs, Geni-

tiv sodann, dem Werwolf, Dativ,

wie man’s nennt, den Werwolf, - damit

hat’s ein End!“ “The banSHEE, in the subject’s place; the banHERS, the possessive case. The banHER, next, is what they call objective case–and that is all.”

“I know that–I’m no men-

tal cripple– The future form and parti-

ciple Are what I crave,” the be-

ast replied. The scholar paused–again

he tried:

Ĉi participoj

plaĉis multe al la fantomo, kaj

rezulte ĝi petis: “Legu plu

el PIV-o pri participoj en

pasivo.”

Dem Werwolf schmei-

chelten die Fälle er rollte seine Au-

genbälle. �Indessen“, bat er,

�füge doch zur Einzahl auch die

Mehrzahl noch!“ The banshee marveled at the cases and writhed with pleasure, making faces, but said: “You did not add, so far, the plural to the singu-

lar!”

…  …  …  …

2/3

“A ‘Will-be-wolf?’ It’s just

too long: ‘Shall-be-wolf?’ ‘Has-been-

wolf?’ Utterly wrong! Such words are wounds

beyond all suture– I’m sorry, but you have no

future.”

La viro ĉagreniĝis tamen: “Jam fermis mi per fin’ kaj amen, ĉar netransiraj bestoj feblas kaj la pasiv’ neniam eblas.”

Der Dorfschullehrer aber musste gestehn, dass er von ihr nichts wusste. Zwar Wölfe gab’s in großer Schar, doch �Wer“ gab’s nur im Sinular.

The teacher, though, admitted then that this was not within his ken. “While bans are frequent”, he advised, “A she cannot be pluri-

zed.”

The Werewolf knew better–his sons still slept

At home, and homewards

now he crept, Happy, humble, without

apology For such folly of philology.

Sed lupfantomo pri transiro ja spertas pli ol eĉ vampiro. La lup’ revenis larmokula al sia famili’ ulula.

Der Wolf erhob sich tränenblind – er hatte ja doch Weib und Kind! Doch da er keine Gelehrter eben, so schied er dankend und ergeben.

The banshee, rising

clammily, wailed: “What about my family?” Then, being not a learned creature, said humbly “Thanks” and left the teacher.

…  …  …  …

3/3

Translation of the German poem “Der Werwolf” by Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) into English by Alexander Gross.

Arg-825-1657 (2013-02-13 18:39:45)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Translation of the German poem “Der Werwolf” by Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) into Esperanto by Bertilo Wennergren.

Arg-825-1653 (2013-02-14 23:29:47)

Vidu  la  re-

tejon http: // bertilow. com/ literaturo/ lupfantomo. html .

Author of this German poem is Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31).

Arg-825-1652 (2013-02-11 18:44:21)

Translation of the German poem “Der Werwolf” by Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) into English by Max Knight (civila nomo: Max Kühnel, ∗1909-06-08 – †1993-08-31).

Arg-825-1660 (2013-02-13 22:36:06)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html . Pri la tradukinto Max Knight vidu: http: // www. jbeilharz. de/ morgenstern/ morgenstern_poems. html kaj http: // www. 123people. de/ ext/ frm? ti= ersonensuche% 20telefonbuch&search_ term= m x% 20k% C3% BChnel&search_ country= DE&st= suc e% 20nach% 20personen&target_ url= aHR0cDovL d3dy5saXRlcmF0dXJoYXVzLmF0L2luZGV4LnBocD9pZD02ODc4 .